Let f(x) = x^2
2 + 6x. Use the definition (ε − δ method) to show that
lim
x→−3
f(x) = −9 .
1
Expert's answer
2020-12-13T16:46:33-0500

We need to prove that for every positive "\epsilon" > 0, there exist a "\delta" such that |f(x) + 9| < "\epsilon" for all x satisfying 0<|x+3|<"\delta\\
|(x^2+6x)+9|<\epsilon\\
|(x+3)(x+3)|<\epsilon\\ \text{for all x satisfying} 0<|x+3|<\delta\\

\text{since |x+3|<} \delta\\
\delta^2<\epsilon\\
\delta<\sqrt\epsilon\\
\text{we choose}\\
\delta=\sqrt\epsilon\\
\text{then the statement}\\
|(x+3)(x+3)|<\epsilon\\ \text{for all x satisfying} 0<|x+3|<\delta\\ \text{holds}\\

Hence\\
Lim_{x\to 2}(x^2+6x)=-9"